1. General status updates/Schedule review.
2. Status of CRV ROC interface.
3. Status of EPICS on DAQ cluster.
4. Status or core ROC firmware testin

CRV data request latency -- 30us of latency (12 + 15 + 1 + 1 + 1) .. Currently there is queueing.
Rick -- already implemented PreFetch request at DTC
Sten -- CRV has to wait for request if no prefetch information because of out of order on-spill requests.
Is it trivial for CRV firmware to allow 50us?

TDAQ Status -
Glenn -- Computer changes have confused things. Daq04 running IOCs. Daq11 running dcs_ots. Need a place for Archiver and mySQL.
Moving EPICS* to mu2edaq12
Bertrand -- working to improve/optimize Tracker modules in Trigger sequence.
Ron -- Verified message facility UDP has file/line number for all severity. Looking into PCIE_* repo to depend on trace_mf.h .. XOAP commander for artdaq ..
Ron updated the kernel on “05 and 06 ” did we do 04? -- always from hwdev. (It was from latest release).
It’s in root directory.
n.b. ELF: I have updated the scripts on all mu2edaq machines to use the version from hwdev.
Rick -- debuggin with CRV ROC -- found DTC bugs, and then back to high rate testing. Experiencing 10-20 us latency even without FEBs.
Eric -- shuffled daq cluster nodes - everyone on daq11. Daq14-17 .. intended for production.
Ryan -- working on artdaq configuration GUI, thinking about how to efficiently instantiate multiple artdaq nodes.
Greg -- Wants to do marker sync tests again. Waiting on Micol.
Vivian -- Waiting on Micol and Monica. Monica is coming tomorrow.
Micol -- Studying Jose’s firmware .. Known Monica-firmware issues not seeing data requests. Testing link communication between DTC-ROC-DTC for reads .. it’s still timing out in software at PCIE.
Glenn/Eric -- we need to switch from SL7 before production. Eric thinks early next year migration complete.